OpenStudy (anonymous):

Calculate the stopping distances for a car that has an initial speed of 95 km/h with a driver's reaction time of 3.0 s for the following values of acceleration.(Reaction time is the time it takes for the driver to initially apply the breaks.) a=-4m/s^2 a=-8m/s^2

OpenStudy (mihai):

a speed of 95km/h in m/s that would be 95000m/3600s=26.38 m/s reaction distance is 3*26.38=79.14m acc=speed/time distance=t*speed, in this case you also add the reaction distance so d=t*s+79.14 -4=26.38/t1 , t1=6.59s, d1=6.59*26.38+79.14=253.11m -8=26.38/t2, t2=3.29s, d2=3.29*26.38+79.13=165.93m